OpenStudy (anonymous):

How do you calculate W if given a negative x force, positive x force and positive y force on a given object with a negative velocity? I am also given mass of 2kg

OpenStudy (anonymous):

Well if what I remember W= F dr where dr is the distance traveled. So first you need to find the force magnitude so the first thing is you want to add the forces in the x coordinate. If the negative force is stronger you should have a force going to the left in the x coordinate and to the right if the positive is stronger. Next you will be left with two forces one in the x coordinate and one in the y coordinate and you want to use Pythagoras to find the magnitude where a is the x force and b is the y force and C will be the magnitude. Once you have the magnitude you find the acceleration due to the force which is F=m (mass) * a (acceleration) where a = F/ m After that you need to know by how much the object traveled and you use the kinematic equation change x = v( t1 - t0) + 1/2 * a * (t1-t0)^2 where v is the velocity and t1 is the final time and t0 is the initial time and a is your acceleration Once you know how far it traveled you use W = F * dr where F= magnitude force dr= change x